OpenStudy (anonymous):

Write the sum using summation notation, assuming the suggested pattern continues. -8 - 3 + 2 + 7 + ... + 67 PLEASEEEE

OpenStudy (anonymous):

@amistre64 @mathstudent55 do u know this by any chance??

OpenStudy (amistre64):

yes

OpenStudy (anonymous):

awesome! how would I go about doing this?

OpenStudy (amistre64):

lets assume since the values are pretty close to one another that the sequence is arithmetic; what defines an arithmetic sequence?

OpenStudy (anonymous):

arithmetic sequence has consecutive terms that are increasing or decreasing at a consistent rate?

OpenStudy (amistre64):

correct, or simply, there is some constant value being added each time -8 + k = -3 -3 + k = 2 2 + k = 7 does this look like itll fit for some k? and if so, what would the value of k be?

OpenStudy (anonymous):

k=5 :D

OpenStudy (amistre64):

good, then lets follow this out a0 = -8 a1 = -8 + 5 a2 = -8 + 5 + 5 a3 = -8 + 5 + 5 + 5 a4 = -8 + 5 + 5 + 5 + 5 an = -8 + 5n , for n=0 to whatever do you agree?

OpenStudy (amistre64):

sometimes its easier to work with when n starts at 1 instead of 0, to adjust for that we just subtract 1 from n an = -8 + 5(n-1) , for n=1 to whatever = -13 + 5n

OpenStudy (anonymous):

ahhhh okay yes I understand that part! now how do I determine if it goes to infinity or just a set number of times like 15?

OpenStudy (amistre64):

well, we know that it has to end at some n such that the rule equals 67: 67 = -13 + 5n; for n=1 to ??

OpenStudy (anonymous):

I meant what I am seeing in my answers is \[\sum_{n=0}^{15}(-8+5n) and \sum_{n=0}^{oo}(-8+5n)\]

OpenStudy (amistre64):

ahh, well we know that n=infinity gets huge .... and -8+5(huge) is way to big to be 67 right? what does -8 +5(15) equal?

OpenStudy (anonymous):

67 so I guess it has to go to infinity right?

OpenStudy (amistre64):

-8 - 3 + 2 + 7 + ... + 67 ^^^ as a rule of thumb, of you have an end to your sequence of numbers .... you just cant have an infinite set.

OpenStudy (amistre64):

it cant go into infinity .... it has a stopping point

OpenStudy (anonymous):

ahh okay so then it stops at 67??

OpenStudy (amistre64):

when n= 0: -8+5(0) = - 8 when n=15: -8+5(15) = 67 so n has to go from 0 to 15 to fit the problem

OpenStudy (amistre64):

yes, it stops at: +67

OpenStudy (anonymous):

that makes sense, so in any arithmetic sequence they can not go to infinity?

OpenStudy (amistre64):

depends on problem spose it just ended it at: -8,-3,2,7,... the three dots at the end there are notation for going on and on and on and on .... to infinity

OpenStudy (amistre64):

but, yours tells us that it starts at -8, goes on a ways following some rule, then stops at 67

OpenStudy (anonymous):

AHHH NOW I UNDERSTAND!!! derp that was a dumb mistake I made

OpenStudy (anonymous):

Thank you so much

OpenStudy (amistre64):

youre welcome